OpenStudy (anonymous):

how do you rationalize the denominator when the denominator is a square root?

OpenStudy (anonymous):

Basically you square something to undo a square-root. It's the same as multiplying the term by itself. \[{{\sqrt{a}}\over{\sqrt{a}}}*{{7}\over{\sqrt{a}}}={{7\sqrt{a}}\over{a}} \]